Let X be a binomial random variable with n = 100 and p = 0.2. Find approximations to these probabilities. (Round your answers to four decimal places.) (c) P(22 < X < 26)
Answer #1
here mean of distribution=μ=np= 20.00
and standard deviation σ=sqrt(np(1-p))= 4.00
for normal distribution z score =(X-μ)/σx
therefore from normal approximation of binomial distribution and continuity correction:
probability =P(22.5<X<25.5)=P((22.5-20)/4)<Z<(25.5-20)/4)=P(0.63<Z<1.38)=0.9162-0.7357=0.1805